# Principal stratification with $U$-statistics under principal ignorability

### v4

@@ Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects have traditionally been the default target of inference, it may not be sufficient when the interest lies in the relative favorability of one potential outcome over the other within the principal stratum. We introduce the principal generalized causal effect estimands to extend the principal average causal effects, accommodating nonlinear contrast functions. Under principal ignorability, we expand the theoretical results in Jiang et al. (2022) to a much broader class of causal estimands in the presence of a binary intermediate variable. We develop identification results and derive the efficient influence functions of the generalized estimands for principal stratification analyses. These efficient influence functions motivate a set of multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on $U$-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
@@ 1 Introduction
The average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the primary estimand of interest in causal inference. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be expanded to a wider class, allowing arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when h(u,v)=uvh(u,v)=u-v, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$. Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$ (see Lu et al. (2020) for an example for bounding such an estimand with ordinal outcomes). However, a closely related quantity, the generalized causal effect, can be formulated:
@@
Two notable examples of the generalized causal effect are the probabilistic index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (Vermeulen et al., 2015), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock et al., 2011). These estimable quantities are alternative measurements for the population effect and carry a causal interpretation (Fay and Li, 2024). In certain cases, such estimands offer advantages that make them relevant in modern applications. First, the probabilistic index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than a unit from $F_{0}$. This formulation generally exhibits greater robustness to distributional features. That is, unlike mean contrasts, which can be distorted by heavy tails, skewness, or outliers, probabilistic indices require fewer distribution assumptions and are less sensitive to irregularities in the outcome distribution (Hauck et al., 2000; Acion et al., 2006). Second, probabilistic indices are also well suited for ordinal outcomes, where the scientific question itself concerns relative ordering rather than pure numerical differences. Hence, as De Schryver and De Neve (2019) noted, “the probabilistic index provides an ordinal answer to an ordinal question”, and it relies only on the rank comparison rather than the metric scale or magnitude of differences. Third, the win-type estimand is now increasingly used for prioritized composite outcomes, where rank-based win-loss comparisons naturally align with clinical priorities across multiple endpoints. Introduced by Pocock et al. (2011), the win ratio has gained substantial traction in clinical trials with composite outcomes, as it offers a principled way to combine endpoints of varying clinical importance while providing an interpretable summary of treatment benefit in terms of the probability of winning versus losing.
@@
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist et al., 1996) and death (Zhang et al., 2009) often have important implications on the definition of a causal effect and require careful adjustment. The principal stratification framework (Frangakis and Rubin, 2002) offers one solution by segmenting the population into principal strata (subpopulations) based on the joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect, referred to as the principal causal effect, can thus be defined and is causally interpretable. However, a predominant focus in prior works has been on the average causal effect within each principal stratum, and there is a lack of related development when the generalized causal effect is of interest within each subpopulation. Therefore, in this article, we propose a new class of estimands, the principal generalized causal effect (PGCE), which accommodates an arbitrary and potentially nonlinear contrast function. As we demonstrate in due course, the PGCE can represent the stratum-specific probabilistic index and strata-specific win ratio with the appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations of interest.
@@
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist et al., 1996) and principal ignorability (Ding and Lu, 2016), which are typically invoked for principal stratification analyses. The monotonicity assumption states that the treatment has a nonnegative causal effect on the intermediate outcome of each unit, restricting the number of potential strata. This condition is considered plausible when, for example, the intermediate variable is death, and the treatment is known to benefit survival, or when the intermediate variable is the actual treatment received and the treatment does not discourage treatment uptake. The principal ignorability assumption requires that the observed pre-treatment covariates are sufficient to control for confounding between the intermediate variable and the outcome. For the principal average causal effect, under monotonicity and principal ignorability, Ding and Lu (2016) formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang et al. (2022) provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and developed multiply robust and locally efficient estimators; Cheng et al. (2023, 2024) extended the multiply robust and locally efficient estimators to estimate principal survival causal effect with a right-censored time-to-event outcome; Tong et al. (2024) expanded multiply robust and locally efficient estimators to estimate fine principal causal effects with multiple treatments. Despite these advancements, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
@@
The contributions of this article are several-fold. First, we propose a class of PGCE estimands in the presence of a binary intermediate variable and develop complementary nonparametric identification formulas under the assumptions of treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and seek efficiency improvement by deriving their corresponding efficient influence functions (Bickel et al., 1993, EIFs). Specifically, we show that when the contrast function is h(u,v)=uvh(u,v)=u-v, our EIFs reduce to those in Jiang et al. (2022) as a special case. More broadly, our framework accommodates pairwise comparison estimands with arbitrary contrast functions under principal stratification, offering a substantially more general treatment and establishing the present work as a standalone contribution. A major complexity for nonlinear contrast functions, however, is that the estimating equations based on the EIFs require nuisance functions defined over pairs of observations, thereby requiring the theory of $U$-statistics to construct feasible estimators. We show that the proposed $U$-statistics are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. When all nuisance functions are correctly specified, our estimators are locally efficient. Finally, the construction of the multiply robust $U$-statistics facilitates the integration of nonparametric estimators for the nuisance functions. To achieve full asymptotic efficiency, we also estimate each nuisance function using data-adaptive machine learners under the cross-fitting scheme (Chernozhukov et al., 2018). We discuss the modifications needed for cross-fitting when a nonlinear contrast function is considered in defining the PGCE, and prove the nonparametric efficiency properties of the debiased machine learning estimators under mild regularity conditions.
The rest of this article is organized as follows. Section 2 introduces our causal estimands and identification assumptions for principal stratification. Section 3 discusses the nonparametric identification formulas for the PGCE and associated moment estimators. Section 4 derives the EIF for each PGCE estimand and develops the EIF-motivated estimators. Section 5 demonstrates the finite-sample performance of the proposed estimators via simulations. Section 6 illustrates the proposed methods with a data example from the United States Jobs Corps study (Schochet et al., 2001), and Section 7 concludes. Detailed proofs of all theoretical results are presented in the Online Supplement.
@@ 2 Notation, estimands and assumptions
2.1 Setting and notation
We consider a study, either experimental or observational, with $n$ individuals and two treatment conditions. We let $Z_{i}$ denote the treatment assignment of individual $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 0 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the Stable Unit Treatment Value Assumption (SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $X_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{X_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d.).
@@ 2.2 Principal generalized causal effects
We operate under the principal stratification framework (Frangakis and Rubin, 2002), and define distinct principal strata according to the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{10,01,00,11\}$. A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
@@
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$ ($S_{1}=s$ and $S_{2}=s$ denote the strata membership for these two selected units). When a linear contrast function is specified as h(u,v)=uvh(u,v)=u-v, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature (Frangakis and Rubin, 2002). Before proceeding, we provide a caveat on the definition of estimands in Remark 1.
@@
One may wonder whether quantities of the form $\tau^{s_{1},s_{2}}_{h}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$ can also be considered. While such cross-stratum estimands are in principle identifiable under structural assumptions, they lack a meaningful causal interpretation. Following the definition of causal effects in Frangakis and Rubin (2002), the comparison of potential outcomes must be made within a well-defined target population. If $s_{1}\neq s_{2}$, then the two potential outcomes $Y_{1}(1)$ and $Y_{2}(0)$ are now drawn from distinct subpopulations characterized by different joint potential intermediate outcomes. In this case, $\tau^{s_{1},s_{2}}_{h}$ represents a cross-world comparison across heterogeneous subgroups rather than a treatment effect within a common subpopulation, and thus does not conform to the spirit of the original principal stratification framework regarding subpopulation-specific causal effects (Frangakis and Rubin, 2002). As an example, when $D$ denotes the survival status in the truncation-by-death setup (Example 1), comparing $Y_{1}(1)$ among the protected (those who survive only under treatment) with $Y_{2}(0)$ among the always-survivors (those who would survive regardless of assignment) conflates two sources of difference: one due to the treatment condition and another due to inherent heterogeneity across subpopulations. This contrast, therefore, cannot be solely attributed to treatment and does not carry a causal interpretation. For this reason, we restrict attention to $\tau_{h}^{s}$ with $s_{1}=s_{2}=s$ to ensure that pairwise comparisons are drawn from the same principal stratum.
Given $S_{1}=S_{2}=s$, we emphasize that the contrast function $h$ is evaluated on two independent replicates drawn from the same principal stratum. This pairwise construction in the causal estimand also motivates the definition of the estimable pairwise mean function in Section 3.1, which serves as the building block for our identification strategy and $U$-statistic estimators. That is, both the estimand and the estimable pairwise mean function require two independent replicates to be properly defined. To illustrate the PGCE estimand, we next present two examples, while leaving the technical details on constructing $U$-statistic estimators to Section 3. Importantly, when a nonlinear contrast function is used to define the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
@@
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is oftentimes only well-defined when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang et al., 2009; Chen et al., 2024). In clinical research, Acion et al. (2006) argued that a more intuitive effect size can be measured by the probabilistic index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probabilistic index (SPI) as
@@
Compared to the SACE, the SPI is less sensitive to the skewness of the distributions of the potential outcomes, heteroskedasticity, and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (Wolfe and Hogg, 1971).
@@
In experimental studies with ordinal outcomes, Wang and Pocock (2016) introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{10,01,00,11\}$ refers to always-takers, compliers, defiers, and never-takers, respectively. Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
@@
The CWR estimand represents the odds of a randomly selected unit from all-treated compliers having a more favorable outcome than one from the all-control compliers subpopulation. Alternatively, the complier win difference (CWD) or complier net benefit estimand is
@@
In the context of noncompliance, Mao (2024) defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE. Mao (2024) employed an instrumental variable approach and discussed the point and interval identification of this estimand, based on monotonicity and the exclusion restriction.

### v5

@@ Abstract
Principal stratification is a popular framework for causal inference in the presence of an intermediate outcome. While the principal average treatment effects are the standard target of inference, they may be insufficient when interest lies in the relative ordering of potential outcomes within a principal stratum. We introduce the principal generalized causal effect estimands to accommodate nonlinear contrast functions, providing robust, probability-scale summaries suitable for ordinal outcomes and win-loss comparisons with composite endpoints. Under principal ignorability, we expand the theoretical results in Jiang2022 to a broader class of causal estimands in the presence of a binary intermediate variable. We develop nonparametric identification results and derive efficient influence functions for the generalized causal estimands in principal stratification analyses. These efficient influence functions motivate multiply robust estimators and lay the ground for obtaining efficient debiased machine learning estimators via cross-fitting based on U-statistics. The proposed methods are illustrated through simulations and the analysis of a data example.
@@ 1 Introduction
The average causal effect—defined as the expected contrast between two potential outcomes under alternative treatment conditions—is often the standard estimand of interest in causal inference. Specifically, let $\{Y(1),Y(0)\}$ denote a pair of potential outcomes for a randomly selected individual from a super-population, and then the average causal effect estimand is defined as $\mathbb{E}\{Y(1)-Y(0)\}$. Conceptually, the average causal effect estimand can be extended to a broader class, enabling arbitrary comparisons. That is, given a contrast function $h(u,v)$ defined on the product space, $\mathcal{Y}\times\mathcal{Y}$ ($\mathcal{Y}$: the potential outcome space), a generalized estimand can be defined as $\mathbb{E}[h\{Y(1),Y(0)\}]$. Examples include the average causal effect when h(u,v)=uvh(u,v)=u-v, and the proportion of the population benefiting from treatment, $\mathbb{P}\{Y(1)>Y(0)\}$, when $h(u,v)=\mathbbm{1}(u>v)$. Because $Y(1)$ and $Y(0)$ are never observed simultaneously, $\mathbb{E}[h\{Y(1),Y(0)\}]$ is generally inestimable for nonlinear $h$ (see lu2020sharp for an example for bounding such an estimand with ordinal outcomes). However, a closely related quantity, the generalized causal effect, can be formulated:
@@
Two notable examples of the generalized causal effect are the probabilistic index (the target parameter for the Mann-Whitney statistic), $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)\}$, when $h(u,v)=\mathbbm{1}(u\geq v)$ (vermeulen2015increasing), and the win ratio, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$, when $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$ (Pocock2012). These estimable quantities are alternative measurements for the population effect and carry a causal interpretation (fay2024causal). In some instances, such estimands offer advantages that make them relevant in social science and biomedical applications. First, the probabilistic index comparing the ranking of outcomes under two potential outcome distributions measures the probability that a randomly chosen unit from $F_{1}$ has a higher-ranked outcome than a unit from $F_{0}$. This formulation generally exhibits greater robustness to distributional features. That is, unlike mean contrasts, which can be distorted by heavy tails, skewness, or outliers, probabilistic indices require fewer distribution assumptions and are less sensitive to irregularities in the outcome distribution (hauck2000generalized; acion2006probabilistic). Second, probabilistic indices are well suited to ordinal outcomes, in which the scientific question concerns relative ordering rather than pure numerical differences. As de2019tutorial noted, “the probabilistic index provides an ordinal answer to an ordinal question”, and it relies only on the rank comparison rather than the metric scale or magnitude of differences. Third, the win-type estimand is now increasingly used for prioritized composite outcomes, where win-loss comparisons naturally align with clinical priorities across multiple endpoints. Introduced by Pocock2012, the win ratio has gained substantial traction in clinical research with composite outcomes, as it offers a principled way to combine endpoints of varying clinical importance while providing an interpretable summary of treatment benefit in terms of the probability of winning versus losing.
@@
When drawing causal inferences from randomized experiments or observational studies, intermediate variables such as treatment noncompliance (Angrist1996) and death (Zhang2008) often have important implications on the definition of a causal effect and require careful adjustment. Principal stratification (Frangakis2002) offers one solution by segmenting the population into principal strata based on the joint potential values of the intermediate variable under alternative conditions. The stratum-specific causal effect, referred to as the principal causal effect, can thus be defined and is causally interpretable. However, prior work has predominantly focused on the average causal effect within each principal stratum, and there has been little related development when the generalized causal effect is of interest within each subpopulation. Therefore, we propose a new class of estimands, the principal generalized causal effect (PGCE), that accommodates an arbitrary, potentially nonlinear contrast function. As we demonstrate in due course, the PGCE can represent the stratum-specific probabilistic index and the stratum-specific win ratio with an appropriate choice of the contrast function, and it can provide a meaningful assessment of the relative treatment effect within subpopulations.
@@
Assuming a binary intermediate variable, we study the identification and inference of the PGCE under two structural assumptions, monotonicity (Angrist1996) and principal ignorability (Ding2016), which are typically invoked for principal stratification analyses. For the principal average causal effect, under monotonicity and principal ignorability, Ding2016 formalized weighting estimators using principal scores (the conditional probabilities of the latent principal strata given observed covariates); Jiang2022 provided alternative identification formulas based on the treatment propensity score, the principal score and the potential outcome models, and developed multiply robust and locally efficient estimators; cheng2023multiply extended the multiply robust and locally efficient estimators to estimating principal survival causal effect with a right-censored time-to-event outcome; tong2024doubly expanded multiply robust and locally efficient estimators to estimate fine principal causal effects with multiple treatments. Despite these advancements, the identification and the semiparametric efficiency theory for the PGCE in the presence of a binary intermediate variable remain largely unclear.
@@
The contributions of this article are several-fold. First, we propose a class of PGCE estimands in the presence of a binary intermediate variable and develop nonparametric identification formulas under the assumptions of treatment unconfoundedness, monotonicity, and principal ignorability. Second, we provide a formal treatment of semiparametric efficiency theory for estimating the PGCE estimands and derive their corresponding efficient influence functions (EIFs, Bickle1993). Specifically, we show that when the contrast function is h(u,v)=uvh(u,v)=u-v, our EIFs reduce to those in Jiang2022 as a special case. More broadly, our framework accommodates pairwise comparison estimands with arbitrary contrast functions, offering a substantially more general treatment and establishing the present work as a standalone contribution. A significant complexity of nonlinear contrast functions, however, is that the estimating equations based on the EIFs require nuisance functions defined on pairs of observations, thereby necessitating the theory of U-statistics to construct feasible estimators. We show that the proposed U-statistics are triply robust in the sense that they are consistent and asymptotically normal when two of the three nuisance functions are correctly specified. When all nuisance functions are correctly specified, our estimators are locally efficient. Finally, the construction of the multiply robust U-statistics facilitates the integration of nonparametric estimators for the nuisance functions. To achieve full asymptotic efficiency, we estimate each nuisance function using data-adaptive machine learners under the cross-fitting scheme (Chernozhukov2018). We discuss the modifications needed for cross-fitting when a nonlinear contrast function is used to define the PGCE, and prove the nonparametric efficiency properties of the debiased machine learning estimators under mild regularity conditions. Example R codes are available at https://github.com/erxc/cusps.
@@ 2 Notation, estimands and assumptions
We consider a study, either experimental or observational, with $n$ individuals and two treatment conditions. We let $Z_{i}$ denote the treatment assignment of individual $i\in\{1,\ldots,n\}$; $Z_{i}=1$ indicates the treatment condition and 0 the control. We use $D_{i}(z)$ and $Y_{i}(z)$ to denote the potential outcomes of the intermediate variable and the final outcome variable of interest for participant $i$ under treatment $Z_{i}=z$ ($z\in\{0,1\}$), respectively. The intermediate variable $D_{i}\in\{0,1\}$ is assumed to be binary, but there is no restriction on the type of variable for $Y_{i}$. Under the Stable Unit Treatment Value Assumption (SUTVA), the observed intermediate and final outcomes can be represented as $D_{i}=Z_{i}D_{i}(Z_{i})+(1-Z_{i})D(1-Z_{i})$ and $Y_{i}=Z_{i}Y_{i}(Z_{i})+(1-Z_{i})Y(1-Z_{i})$. Let $\bm{X}_{i}$ represent the vector of observed pre-treatment covariates, and the observed data for participant $i$ is thus denoted by $\mathcal{O}_{i}=\{\bm{X}_{i},Z_{i},D_{i},Y_{i}\}$, where $\mathcal{O}_{i}$’s are assumed to be independent and identically distributed (i.i.d.).
@@ 2.1 Principal generalized causal effects
We operate within the principal stratification framework (Frangakis2002) and define distinct principal strata based on the joint potential values of the intermediate variable under alternative conditions. Specifically, we use $S=\{D(1),D(0)\}$ to denote the principal stratum variable, which has four potential values: $\{10,01,00,11\}$. A key insight under the principal stratification framework is that although $D$ is a post-randomization variable, the stratum variable $S$ is unaffected by the assignment and can be viewed as a pre-treatment variable. Therefore, based on any contrast function $h$, we define the principal generalized causal effect (PGCE) estimand as
@@
where $Y_{1}(1)$ and $Y_{2}(0)$ denote the potential outcomes for two randomly selected units from the subpopulation with strata membership $s$. When a linear contrast function is specified as h(u,v)=uvh(u,v)=u-v, the PGCE in (2) reduces to $\tau_{h}^{s}=\mathbb{E}\{Y(1)-Y(0)|S=s\}$, which is precisely the principal average causal effect studied in the existing principal stratification literature (Frangakis2002).
@@
One may wonder whether quantities of the form $\tau^{s_{1},s_{2}}_{h}=\mathbb{E}[h\{Y_{1}(1),Y_{2}(0)\}|S_{1}=s_{1},S_{2}=s_{2}]$ with $s_{1}\neq s_{2}$ can also be considered. While such cross-stratum estimands are in principle identifiable under structural assumptions, they lack a meaningful causal interpretation. Following the definition of causal effects in Frangakis2002, the comparison of potential outcomes must be made within a well-defined target population. If $s_{1}\neq s_{2}$, then the two potential outcomes $Y_{1}(1)$ and $Y_{2}(0)$ are now drawn from distinct subpopulations characterized by different joint potential intermediate outcomes. In this case, $\tau^{s_{1},s_{2}}_{h}$ represents a cross-world comparison across heterogeneous subgroups rather than a treatment effect within a common subpopulation, and thus does not conform to the spirit of the original principal stratification framework regarding subpopulation-specific causal effects (Frangakis2002). As an example, when $D$ denotes the survival status in the truncation-by-death setup (Example 1), comparing $Y_{1}(1)$ among the protected (those who survive only under treatment) with $Y_{2}(0)$ among the always-survivors (those who would survive regardless of assignment) conflates two sources of difference: one due to the treatment condition and another due to inherent heterogeneity across subpopulations. This contrast, therefore, cannot be solely attributed to treatment and does not carry a straightforward causal interpretation. For this reason, we restrict attention to $\tau_{h}^{s}$ with $s_{1}=s_{2}=s$.
Given $S_{1}=S_{2}=s$, we emphasize that the contrast function $h$ is evaluated on two independent replicates drawn from the same principal stratum. This pairwise construction in the causal estimand also motivates the definition of the estimable pairwise mean function in Section 3.1, which serves as the building block for our identification strategy and U-statistic-based estimators. To illustrate the PGCE estimand, we next present two examples, while leaving the technical details on constructing U-statistic-based estimators to Section 3. Importantly, when a nonlinear contrast function is used to define the PGCE, $\tau_{h}^{s}$ is no longer a simple average of $Y(z)$ within each principal stratum.
@@
When $D$ represents the survival status (alive if $D=1$) that is ascertained before the measurement of the final outcome $Y$ (such as quality of life, which is only well-defined when $D=1$), there exist four principal strata based on the potential survival status. That is, $s=11$ represents the always-survivors who would survive until the measurement of the final outcome regardless of the assignment; $s=10$ and $s=01$ represent the protected and harmed, who would survive until the measurement of the final outcome only under the treatment condition and control condition, respectively; $s=00$ represents the never-survivors who would not survive until the measurement of the final outcome regardless of the assignment. Because the final outcome is typically only well-defined when $D=1$, the survivor average causal effect (SACE) has been suggested as a relevant target of inference with $\mathrm{SACE}=\mathbb{E}\{Y(1)-Y(0)|S=11\}$ (Zhang2008; Chen2024). In clinical research, acion2006probabilistic argued that a more intuitive effect size can be measured by the probabilistic index defined with a contrast function $h(u,v)=\mathbbm{1}(u\geq v)$. Based on this contrast function, we can define the survivor probabilistic index (SPI) as
@@
Compared with the SACE, the SPI is less sensitive to skewness in the distributions of potential outcomes, heteroskedasticity, and outliers. In the special case of normality and under the equal variance assumption, the SPI is a strictly monotone function of the SACE because $\mathrm{SPI}=\Phi(\mathrm{SACE}/\sqrt{2}\sigma)$, where $\Phi$ is the standard Gaussian cumulative distribution function and $\sigma^{2}$ is the constant variance of the potential outcomes (wolfe1971constructing).
@@
In experimental studies with ordinal outcomes, wang2016win introduced the win ratio estimand for pairwise comparisons, $\mathbb{P}\{Y_{1}(1)>Y_{2}(0)\}/\mathbb{P}\{Y_{1}(1)<Y_{2}(0)\}$. This estimand corresponds to a two-dimensional contrast function specified as $h(u,v)=\{\mathbbm{1}(u>v),\mathbbm{1}(u<v)\}$. If $D$ represents the actual treatment received that can differ from the assignment, then $S\in\{10,01,00,11\}$ refers to always-takers, compliers, defiers, and never-takers, respectively. Therefore, the win ratio estimands can be extended to define the compliers win ratio (CWR) as
@@
The CWR represents the odds that a randomly selected complier under the treatment condition has a more favorable outcome than under the control condition. Alternatively, the complier win difference (CWD) or complier net benefit estimand is
@@
In the context of noncompliance, Mao2024 defined a local Mann-Whitney treatment effect estimand given by $\mathbb{P}\{Y_{1}(1)\geq Y_{2}(0)|S_{1}=S_{2}=10\}$, which is also a special case of the PGCE. Mao2024 employed an instrumental variable approach and discussed the point and interval identification of this estimand under monotonicity and exclusion restriction.